OpenStudy (anonymous):

since i am new to the GCSE system, can anybody explain what is the Sine rule and the Cosine rule? I did not learn it in my previous school in grade 10th. :(

OpenStudy (anonymous):

sine rule as in the law of sines?

OpenStudy (saifoo.khan):

Signs on Roads.

OpenStudy (anonymous):

yes... satellite73 and saifoo.khan, you always joke!

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=oQwszMUiOws

OpenStudy (saifoo.khan):

im sure you can use youtube.

OpenStudy (anonymous):

says \[\frac{\sin(\alpha)}{a}=\frac{\sin(\beta)}{b}=\frac{\sin(\gamma)}{c}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

i draw like a two year old

OpenStudy (anonymous):

go here http://justmathtutoring.com/ go to "trigonometry" and look at "law of sines" and "law of cosines"